# And yet it moves: why irrotational waves induce a mean Lagrangian drift

### v2

@@ 3.2 Drift and Energy
is conserved following the flow (i.e. the material derivative of $\Gamma$ is zero). A less commonly known Lagrangian version of this theorem (see Salmon, 1988) equivalently defines the circulation as
@@
$\frac{{\rm d}\Gamma}{{\rm d}t}=\oint\frac{\partial\bm{A}}{\partial t}\bm{\cdot}{\rm d}\bm{\alpha}=0\,,$ (18)
@@
where we used Stokes theorem and the fact that the curl of a gradient always vanishes. Additionally, and importantly, if we constrain ourselves to periodic flows, such as those represented by (12), then we can choose a closed contour as in figure 1 which is a rectangle in label space with width $\lambda=2\pi/k$, extending vertically from $(\beta=\beta_{0})$ to the infinite bottom $(\beta\rightarrow-\infty)$ where the velocity vanishes. Because the domain is periodic, the contributions from the sides cancel out, and due to our deep water condition the bottom boundary does not contribute. Thus, (19) reduces to
Figure 1: Schematic of a potential closed material loop for periodic, progressive waves (red). The top contour is a material line of constant vertical label $\beta=\beta_{0}$. Because the entire domain is $\lambda$-periodic, the side contours cancel. Due to our infinite bottom condition, $\bm{A}$ vanishes as we approach the bottom and there are no contributions there. Note the clockwise orientation used.
@@
or equivalently, that the phase average of $\dot{x}x_{\alpha}+\dot{y}y_{\alpha}$ is zero for any irrotational and horizontally periodic fluid. From here we can take advantage of our expansions (12) which relate $\alpha$ and $t$ derivatives as
@@
where $T=\tfrac{1}{2}(\dot{x}^{2}+\dot{y}^{2})$ is defined as the kinetic energy density. This novel relation implies that the mean Lagrangian drift is linked to the kinetic energy of the system through the phase speed $c$, with all of its nonlinear corrections, implying that any irrotational wave of the form (12) that has energy must also have a drift, even if the underlying system does not represent surface gravity waves. Recall that all we have invoked here is Kelvin’s circulation theorem, irrotational flow, and trajectories of the form (12).
Lastly, we emphasize that this relationship holds level-wise (i.e. for each vertical $\beta$ level) and as such encodes depth dependence. A similar relationship linking momentum density to kinetic energy density in the Eulerian frame was first found by Levi-Civita (1924), written in the form
@@
This was later found to hold between any two material contours, or equivalently streamlines in a co-moving frame, by Starr (1947). While this equation is similar in scope to (25), it only holds when vertically integrated between material contours, which are more naturally formulated in the Lagrangian frame as lines of constant $\beta$. The relationship found above (25) holds for each material line of constant $\beta$, and as such shows equivalent vertical dependence in $U(\beta)$ and $\langle T\rangle$, but it is also more connected to the classical meanings of terms such as momentum and kinetic energy densities, which in the Lagrangian frame directly encodes the mean Lagrangian drift $U(\beta)$.
@@ 3.3 Drift, Mean Water Level, and Mean Pressure
$\text{MWL}_{\text{Lag}}$ $=\frac{1}{\lambda}\int_{0}^{\lambda}y(\alpha,0,t)\,{\rm d}\alpha=y_{0}(b)$ (30)
@@
and thus the mean water levels are not the same. Precisely, they differ within the integral by the factor $x_{\alpha}$, which corresponds to the unequal spacing of particles along the water surface. In surface gravity waves this tends to bunch particles towards the wave crest and spread them out within the trough. A consequence of this is that the wave crests and troughs, as defined by Lagrangian phase in (12) are also of unequal lengths in physical space. The incompressibility condition in the Lagrangian frame is maintained by stretching or compressing of particles in the vertical.
The precise value for the Lagrangian mean water level $y_{0}(\beta)$ is gauge dependent, though the gauge we choose for irrotational water waves is $\mathcal{J}=1$, which just means that areas in label space map equivalently to areas in physical space. We can then constrain its value by the enforcement that the Eulerian mean water level is zero. Thus we choose $y_{0}(\beta)$ such that
@@
$p+gy+\frac{1}{2}(\dot{x}^{2}+\dot{y}^{2})-c\dot{x}+\frac{c^{2}}{2}=f(\alpha,t)\,,$ (38)
where $f(\alpha,t)$ is a constant of integration, whose value is constrained by the boundary conditions. For our system, as we approach the infinite bottom, all wave terms $(\dot{x},\dot{y},U,y_{0})$ vanish, and pressure becomes hydrostatic $p\rightarrow-g\beta$, which implies $f(\alpha,t)=\frac{c^{2}}{2}$. Clearly, this looks like Bernoulli’s equation in Lagrangian coordinates as was first noticed by Pizzo et al. (2023). If we take the phase average of this equation, we find
@@
where $\langle V\rangle=g(\beta+y_{0}(\beta))$ is the average potential energy of particles. Thus the mean pressure acts as a Lagrangian for the system, which is similar to what Luke (1967) found in the Eulerian frame. If we apply Whitham’s method using the averaged Lagrangian, and substitute the expansions (12) for $\langle T\rangle$ and $\langle V\rangle$, the action becomes
$\mathcal{A}=\int_{t_{1}}^{t_{2}}\int_{-\infty}^{0}\frac{1}{2}U(\beta)^{2}+\frac{1}{4}\sum_{n}n^{2}k^{2}(x_{n}+y_{n})^{2}(c-U(\beta))^{2}-gy_{0}(\beta)\,{\rm d}\beta\,{\rm d}t\,.$ (42)
@@
exactly the same as (13) which was originally found by a purely kinematic argument.
@@ 4 Conservation Laws
The previous section introduced a close connection between momentum and energy for periodic, irrotational waves in a fluid. What are the corresponding conservation laws for these quantities? Returning to the momentum equations
@@
$\frac{\partial}{\partial t}\underbrace{\int_{-\infty}^{0}\mathcal{J}\dot{x}\,{\rm d}\beta}_{\equiv\mathcal{I}}+\frac{\partial}{\partial\alpha}\underbrace{\int_{-\infty}^{0}py_{\beta}\,{\rm d}\beta}_{\equiv S}=py_{\alpha}\Big{|}_{b=0}\,,$ (50)
@@
$\frac{\partial\mathcal{I}}{\partial t}+\frac{\partial S}{\partial\alpha}=py_{\alpha}\Big{|}_{\beta=0}\,,$ (51)
@@
$\frac{\partial E}{\partial t}+\frac{\partial F}{\partial\alpha}=p(\dot{x}y_{\alpha}-x_{\alpha}\dot{y})\Big{|}_{\beta=0}\,,$ (53)
@@
$\frac{\partial\langle\mathcal{I}\rangle}{\partial t}=0\,,$ (56)
@@ 4.1 Generating Stokes waves from rest
While the previous analysis showed why a drift must occur if the wave is to be progressive, irrotational and contain energy, it is helpful to also show how a mean Lagrangian drift can arise on an initially quiescent flow. To begin, we consider a still fluid which at $t=0$ is subject to an external wavelike surface pressure forcing (e.g. by wind)
$p(\beta=0)=\epsilon p_{0}\sin(k\alpha-\omega t)\,,$ (57)
where $\epsilon\ll 1$ is our small parameter, and we take $\omega=\sqrt{gk}$ so that the pressure disturbance propagates at the same speed as a surface gravity wave with the same wavelength. Physically speaking, this pressure forcing drives a resonant response in the sea surface, generating waves whose amplitudes grow linearly in time. In the Lagrangian frame, the particle trajectories for this system valid to second order in $\epsilon$ are found to be
$x(\alpha,\beta,t)$ $=\alpha+\frac{\epsilon\omega p_{0}}{2g}te^{k\beta}\sin(k\alpha-\omega t)+\frac{\epsilon^{2}p_{0}^{2}k^{2}\omega}{12g}e^{2k\beta}t^{3}\,,$ (58)
$y(\alpha,\beta,t)$ $=\beta-\frac{\epsilon\omega p_{0}}{2g}te^{k\beta}\cos(k\alpha-\omega t)+\frac{\epsilon^{2}p_{0}^{2}k^{2}}{8g}e^{2k\beta}t^{2}\,,$ (59)
$p(\alpha,\beta,t)$ $=-g\beta++\epsilon p_{0}e^{k\beta}\sin(k\alpha-\omega t)-\frac{\epsilon^{2}p_{0}^{2}k}{8g}(e^{2k\beta}-1)\,.$ (60)
Note that we have ignored the mean Lagrangian drift in the phase since it does not affect the results to second order. Because the amplitude grows linearly in time, the mean Lagrangian drift which is normally proportional to the square of the amplitudes times $t$ correspondingly grows as $t^{3}$. The second order term in (59) is just the mean water level which scales as the square of the amplitude, which is there to ensure the incompressibility condition subject to the gauge $\mathcal{J}=1$. To see how this explicitly connects to the drift, we first will perform a phase average of the horizontal momentum conservation law (51)
$\frac{\partial}{\partial t}\int_{-\infty}^{0}\langle\dot{x}\rangle\,{\rm d}\beta=\frac{\partial}{\partial t}\int_{-\infty}^{0}U(\beta,t)\,{\rm d}\beta=\langle py_{\alpha}\big{|}_{\beta=0}\rangle\,.$ (61)
This gets rid of the flux terms since the entire solution is periodic in $\alpha$. Our gauge choice $\mathcal{J}=1$ trivializes the Jacobian term. Thus the mean external pressure forcing, represented by the correlation of $p$ and $y_{\alpha}$ at the surface, provides a source of horizontal momentum which fuels the increase of the mean horizontal momentum, or equivalently the vertically integrated Lagrangian mean drift.
@@
$\frac{\partial}{\partial t}\int_{-\infty}^{0}U(\beta,t)\,{\rm d}\beta=\langle py_{\alpha}\big{|}_{\beta=0}\rangle=\frac{\epsilon^{2}p_{0}^{2}\omega k}{4g}t\,,$ (62)
@@
We can do the same analysis for the wave energy, multiplying (46) by $\dot{x}$ and (47) by $\dot{y}$, adding them together and vertically integrating to get
$\frac{\partial}{\partial\tau}\int_{-\infty}^{0}\mathcal{J}\bigg{(}\frac{\dot{x}^{2}+\dot{y}^{2}}{2}+gy\bigg{)}\,{\rm d}\beta+\frac{\partial}{\partial\alpha}\int_{-\infty}^{0}p(x_{\beta}\dot{y}-\dot{x}y_{\beta})\,{\rm d}\beta=p(\dot{x}y_{\alpha}-x_{\alpha}\dot{y})\Big{|}_{\beta=0}\,,$ (63)
where we define
as the energy density (kinetic plus potential) and
$F\equiv\int_{-\infty}^{0}p(x_{\beta}\dot{y}-\dot{x}y_{\beta})\,{\rm d}\beta\,,$ (65)
as the energy flux. When pressure is zero at the surface, total energy is conserved. Going back to our simple example, we perform a phase average to recover
$\frac{\partial\langle E\rangle}{\partial t}=p(\dot{x}y_{\alpha}-x_{\alpha}\dot{y})\Big{|}_{\beta=0}=\frac{1}{4}\epsilon^{2}kp_{0}^{2}t+O(\epsilon^{4})\,,$ (66)
which shows that, just as with momentum, generating waves requires a flux of energy from the wind to the waves. However to lowest order, this energy resides wholly in the orbital particle motion and gravitational potential. Recalling the result from the previous section, we do indeed see that the kinetic energy and mean Lagrangian drift are related, notably that the mean source of momentum multiplied by $c=\sqrt{g/k}$ is equivalent to the mean source of energy at this order, once again highlighting the connection between these two quantities.
@@ 5 Discussion
In this paper we showed that the mean Lagrangian drift, equivalent to the phase averaged momentum density in the physically-motivated Lagrangian frame, is intimately connected to the vorticity and energy densities for irrotational, monochromatic and periodic waves. We further highlighed this connection by showing that the sources of momentum and energy from the perturbation all add to the momentum and energy of the wave field using a simple example of an initially quiescent fluid resonantly forced by a wavelike pressure disturbance at the surface. Physically speaking, this implies that the wave induced mean Lagrangian drift arises due to the necessary input momentum and energy to generate an irrotational wave from rest by any conservative force.
@@
Our novel theoretical results imply that for irrotational, monochromatic and periodic waves, the mean kinetic energy and momentum of particles are intimately related through the wave’s phase speed. This could suggest that anywhere energy is jettisoned, such as by wave breaking, it is accompanied by a shedding of momentum to the underlying mean flow (Rapp & Melville, 1990). This connection between wave energy and mean momentum helps to illuminate the close two-way coupling between currents and waves, and as such will be of particular interest to the air-sea interaction community seeking to model transport and energy budgets between the atmosphere and ocean.
@@
In addition, we explored a connection between the mean potential energy, mean kinetic energy, and mean pressure, showing that at the surface, mean kinetic and mean potential energies are equal, which relates the surface drift directly to the mean water level. This is especially relevant to the observational community as direct measurements of the mean Lagrangian drift are particularly difficult (Kenyon, 1969), especially close to the surface (Lenain & Pizzo, 2020). On the other hand, measurements of geometric properties such the Lagrangian mean water level, might offer an alternative way to estimating the mean Lagrangian drift, though the sensitivity and accuracy of this method is under future consideration by the authors.
@@
It should be noted that the results presented here are similar to the “pseudomomentum rule” in Generalized Lagrangian Mean (GLM) theory, which states that $O(A^{2})$ mean forces can be calculated as if pseudomomentum were momentum, and the fluid medium were absent (McIntyre, 2019), where $A$ is the small wave amplitude. For surface waves, pseudomomentum per unit mass is defined as the wave energy over the phase speed $c$. We emphasize that it is twice the mean kinetic energy, which when divided by $c$, yields $U(\beta)$, valid to all orders of amplitude. To lowest order, the mean kinetic and potential energies are equal, which explains the $O(A^{2})$ result. The exact difference between mean kinetic and potential energy is given by the mean pressure (41), which has nonvanishing terms starting at $O(A^{4})$.
@@
Note that these results do not hold in the general case of rotational waves, such as Gerstner waves, which have no mean Lagrangian drift and therefore no net momentum density, yet still have non-zero energy from their orbital motion. In this case, the connection between kinetic energy and drift fails due to the non-vanishing circulation. Here we focused solely on irrotational flow, though if desired any arbitrary vorticity could be prescribed to the system, generating a nonzero circulation which would balance the kinetic energy term in lieu of the drift. A comprehensive investigation of rotational flow is explored in the appendix.

### v3

@@ 3.2 Drift and Energy
is conserved following the flow (i.e. $\dot{\Gamma}=0$). A lesser known Lagrangian version of this theorem (see Salmon, 1988, eq. 4.12) equivalently defines the circulation as
@@
$\frac{\partial\Gamma}{\partial\tau}=\oint\frac{\partial\bm{A}}{\partial\tau}\bm{\cdot}{\rm d}\bm{\alpha}=0\,,$ (20)
@@
where we used Stokes theorem and the fact that the curl of a gradient always vanishes. Additionally, and importantly, if we constrain ourselves to spatially periodic flows, such as those represented by (13), then we can choose a closed contour as in figure 1 which is a rectangle in label space with width $\lambda=2\pi/k$, extending vertically from $(\beta=\beta_{0})$ to the infinite bottom $(\beta\rightarrow-\infty)$ where the velocity vanishes. Because $\vec{A}$ is $\lambda$-periodic, which can be shown from (13), the side contours cancel, and due to our deep water condition the bottom boundary does not contribute. Therefore the only segment in the contour which contributes is the top segment, reducing (21) to
Figure 1: Schematic of a potential closed material loop for periodic, progressive waves (red). The top contour is a material line of constant vertical label $\beta=\beta_{0}$. Because $\vec{A}$ is $\lambda$-periodic, which can be shown from (13), the side contours cancel. Due to our infinite bottom condition, $\bm{A}$ vanishes as we approach the bottom and there are no contributions there. Note the clockwise orientation used.
@@
or equivalently, that the phase average of $\dot{x}x_{\alpha}+\dot{y}y_{\alpha}$ is zero for any irrotational and horizontally periodic fluid. From here we can take advantage of our expansions (13) which relate $\alpha$ and $\tau$ derivatives as
@@
where $T=\tfrac{1}{2}(\dot{x}^{2}+\dot{y}^{2})$ is defined as the kinetic energy density. This relation implies that the mean Lagrangian drift is linked to the kinetic energy of the system through the phase speed $c$, with all of its nonlinear corrections, implying that any irrotational wave of the form (13) that has energy must also have a mean Lagrangian flow, even if the underlying system does not represent surface gravity waves. Recall that all we have invoked here is Kelvin’s circulation theorem, irrotational flow, and trajectories of the form (13). Lastly, we emphasize that this relationship holds level-wise (i.e. for each vertical $\beta$ level) and as such encodes depth dependence. A similar relationship linking momentum density to kinetic energy density in the Eulerian frame was first found by Levi-Civita (1924), written in the form
@@
This was later found to hold between any two material contours, or equivalently streamlines in a co-moving frame, by Starr (1947). While this equation is similar in scope to (27),that there is a direct connection between momentum and kinetic energy, these terms mean different things in different frames. To see this, consider (29), the Eulerian averaged Eulerian momentum over one period. One can split this integral at the still water level $y=0$ yielding
$I=\overline{\int_{-h}^{0}u\,{\rm d}y}+\overline{\int_{0}^{\eta(x,t)}u\,{\rm d}y}\,.$ (31)
Since the first integral’s endpoints are constants, we can move the Eulerian phase average inside the integral. But since $\overline{u}=0$ for irrotational surface gravity waves, the first integral is identically zero. Thus the only net Eulerian momentum density exists in the second integral between the still water level and the sea surface, or in other words, between the crests and troughs. Geometrically speaking, this occurs because at physical locations between the crests and troughs, a fixed $(x,y)$ point is outside of the fluid part of the time and thus does not experience a full period within the fluid. Putting aside the difficulty of dealing with points partly outside of the fluid domain, these conserved Eulerian quantities are not physically connected to the mass flux of particles, and one would not be able to isolate the mean Lagrangian drift from such an approach. The relationship found above (27) holds for each material line of constant $\beta$, and as such shows equivalent vertical dependence in $U(\beta)$ and $\langle T\rangle$, but it is also more connected to the classical meanings of terms such as momentum and kinetic energy densities, which in the Lagrangian frame directly encodes the mean Lagrangian drift $U(\beta)$.
@@ 3.3 Drift, Mean Water Level, and Mean Pressure
$\text{MWL}_{\text{Lag}}$ $=\frac{1}{\lambda}\int_{0}^{\lambda}y(\alpha,0,t)\,{\rm d}\alpha=y_{0}(\beta)$ (33)
@@
and thus the mean water levels are not the same. Precisely, they differ within the integral by the factor $x_{\alpha}$, which corresponds to the unequal spacing of particles along the water surface. In surface gravity waves this tends to bunch particles towards the wave crest and spread them out within the trough. A consequence of this is that the wave crests and troughs, as defined by Lagrangian phase in (13) are also of unequal lengths in physical space. The incompressibility condition in the Lagrangian frame is therefore maintained by stretching or compressing of particles in the vertical. We choose $y_{0}(\beta)$ such that the mean water level is zero, that is
@@
$p+gy+\frac{1}{2}(\dot{x}^{2}+\dot{y}^{2})-c\dot{x}+\frac{c^{2}}{2}=f(\alpha,\tau)\,,$ (41)
where $f(\alpha,\tau)$ is a constant of integration, whose value is constrained by the boundary conditions. For our system, as we approach the infinite bottom, all wave terms $(\dot{x},\dot{y},U,y_{0})$ vanish, and pressure becomes hydrostatic $p\rightarrow-g\beta$, which implies $f(\alpha,\tau)=c^{2}/2$. Clearly, this looks like Bernoulli’s equation in Lagrangian coordinates as was first noticed by Pizzo et al. (2023). If we take the phase average of this equation, we find
@@
where $\langle V\rangle=g(\beta+y_{0}(\beta))$ is the average potential energy of particles. Thus the mean pressure acts as a Lagrangian for the system, which is similar to what Luke (1967) found in the Eulerian frame, where the expression of the pressure from Bernoulli’s equation was used as a Lagrangian to get both the equations of motion and boundary conditions for the velocity potential $\phi$ and surface $\eta$. If we apply Whitham’s method using the averaged Lagrangian, and substitute the expansions (13) for $\langle T\rangle$ and $\langle V\rangle$, the action becomes
$\mathcal{A}=\int_{t_{1}}^{t_{2}}\int_{-\infty}^{0}\frac{1}{2}U(\beta)^{2}+\frac{1}{4}\sum_{n}n^{2}k^{2}(x_{n}+y_{n})^{2}(c-U(\beta))^{2}-gy_{0}(\beta)\,{\rm d}\beta\,{\rm d}\tau\,.$ (45)
@@
exactly the same as (15) which was originally found by a dynamic constraint.
@@ 4 Conservation Laws
The previous section introduced a close connection between momentum and energy for spatially periodic, irrotational waves in a fluid. What are the corresponding conservation laws for these quantities? Returning to the momentum equations
@@
$\frac{\partial}{\partial\tau}\underbrace{\int_{-\infty}^{0}\mathcal{J}\dot{x}\,{\rm d}\beta}_{\equiv\mathcal{I}}+\frac{\partial}{\partial\alpha}\underbrace{\int_{-\infty}^{0}py_{\beta}\,{\rm d}\beta}_{\equiv S}=py_{\alpha}\Big{|}_{b=0}\,,$ (53)
@@
$\frac{\partial\mathcal{I}}{\partial\tau}+\frac{\partial S}{\partial\alpha}=py_{\alpha}\Big{|}_{\beta=0}\,,$ (54)
@@
$\frac{\partial E}{\partial\tau}+\frac{\partial F}{\partial\alpha}=p(\dot{x}y_{\alpha}-x_{\alpha}\dot{y})\Big{|}_{\beta=0}\,,$ (56)
@@
$\frac{\partial\langle\mathcal{I}\rangle}{\partial\tau}=0\,,$ (59)
@@ 4.1 Generating Stokes waves from rest
While the previous analysis showed why a drift must occur if the wave is progressive, irrotational and contains energy, it is helpful to also show how a mean Lagrangian drift can arise on an initially quiescent flow. To begin, we consider a still fluid which at $\tau=0$ is subject to an external wavelike surface pressure forcing (e.g. by wind)
$p(\beta=0)=\epsilon p_{0}\sin(k\alpha-\omega\tau)\,,$ (60)
where $\epsilon\ll 1$ is our small parameter, and we take $\omega=\sqrt{gk}$ so that the pressure disturbance propagates at the same speed as a surface gravity wave with the same wavelength. Physically speaking, this pressure forcing drives a resonant response in the sea surface, generating waves whose amplitudes grow linearly in time (Pizzo & Wagner, 2021). In the Lagrangian frame, the particle trajectories for this system valid to second order in $\epsilon$ are found to be
$x(\alpha,\beta,\tau)$ $=\alpha+\frac{\epsilon\omega p_{0}}{2g}\tau e^{k\beta}\sin(k\alpha-\omega\tau)+\frac{\epsilon^{2}p_{0}^{2}k^{2}\omega}{12g}e^{2k\beta}\tau^{3}\,,$ (61)
$y(\alpha,\beta,\tau)$ $=\beta-\frac{\epsilon\omega p_{0}}{2g}\tau e^{k\beta}\cos(k\alpha-\omega\tau)+\frac{\epsilon^{2}p_{0}^{2}k^{2}}{8g}e^{2k\beta}\tau^{2}\,,$ (62)
$p(\alpha,\beta,\tau)$ $=-g\beta++\epsilon p_{0}e^{k\beta}\sin(k\alpha-\omega\tau)-\frac{\epsilon^{2}p_{0}^{2}k}{8g}(e^{2k\beta}-1)\,,$ (63)
where solutions are found through a standard perturbation approach (see Salmon, 2020, ch 1. for an outline of the method). Note that we have ignored the mean Lagrangian drift in the phase since it does not affect the results to second order. Because the amplitude grows linearly in time, the mean Lagrangian drift which is normally proportional to the square of the amplitudes times $\tau$ correspondingly grows as $\tau^{3}$. The second order term in (62) is just the mean water level which scales as the square of the amplitude, which is there to ensure the incompressibility condition subject to the gauge $\mathcal{J}=1$. To see how this explicitly connects to the drift, we first will perform a phase average of the horizontal momentum conservation law (54)
$\frac{\partial}{\partial\tau}\int_{-\infty}^{0}\langle\dot{x}\rangle\,{\rm d}\beta=\frac{\partial}{\partial\tau}\int_{-\infty}^{0}U(\beta,\tau)\,{\rm d}\beta=\langle py_{\alpha}\big{|}_{\beta=0}\rangle\,.$ (64)
This gets rid of the flux terms since the entire solution is spatially periodic in $\alpha$. Our gauge choice $\mathcal{J}=1$ trivializes the Jacobian term. Thus the mean external pressure forcing, represented by the correlation of $p$ and $y_{\alpha}$ at the surface, provides a source of horizontal momentum which fuels the increase of the mean horizontal momentum, or equivalently the vertically integrated Lagrangian mean drift.
@@
$\frac{\partial}{\partial\tau}\int_{-\infty}^{0}U(\beta,\tau)\,{\rm d}\beta=\langle py_{\alpha}\big{|}_{\beta=0}\rangle=\frac{\epsilon^{2}p_{0}^{2}\omega k}{4g}\tau\,,$ (65)
@@
As a further check, we see by inserting our solutions (61) – (63) into the phase average of the energy conservation law (56), we recover
$\frac{\partial\langle E\rangle}{\partial\tau}=\Big{\langle}p(\dot{x}y_{\alpha}-x_{\alpha}\dot{y})\Big{|}_{\beta=0}\Big{\rangle}=\frac{1}{4}\epsilon^{2}kp_{0}^{2}\tau+O(\epsilon^{4})\,,$ (66)
which shows that, just as with momentum, generating waves requires a flux of energy from the wind to the waves. However to lowest order, this energy resides wholly in the orbital particle motion and gravitational potential energy from the mean water level. Recalling the result from the previous section, we do indeed see that the kinetic energy and mean Lagrangian drift are related, notably that the mean source of momentum multiplied by $c=\sqrt{g/k}$ is equivalent to the mean source of energy at this order, once again highlighting the connection between these two quantities.
@@ 5 Discussion
In this paper we showed that the mean Lagrangian drift, equivalent to the phase averaged momentum density in the physically-motivated Lagrangian frame, is intimately connected to the vorticity and energy densities for irrotational, monochromatic and spatially periodic waves. We further highlighed this connection by showing that sources of momentum and energy (e.g. from the wind) all add to the momentum and energy of the wave field using a simple example of an initially quiescent fluid resonantly forced by a wavelike pressure disturbance at the surface. Physically speaking, this implies that the wave induced mean Lagrangian drift arises due to the necessary input momentum and energy to generate an irrotational wave from rest by any conservative force. Thus we are well equipped to answer the questions posed in the introduction. Permanent, progressive and irrotational waves require a mean motion of water due to the fact that for these waves to have kinetic energy, they require a net momentum (or mass flux), which in the Lagrangian frame resides in the mean Lagrangian drift. Its magnitude and direction are set by the strict dynamic constraint of irrotational flow, as prescribing the vorticity on particles is equivalent to prescribing their mean Lagrangian drift (see the appendix for cases with nonvanishing vorticity). Finally, the mean Lagrangian drift is not simply related to the mean kinetic energy density in a bulk sense, it is exactly proportional to it, at all vertical material levels, with a factor of $1/c$.
@@
Our theoretical results imply that for irrotational, monochromatic and periodic waves, the mean kinetic energy and momentum of particles are intimately related through the wave’s phase speed. This could suggest that anywhere energy is jettisoned, such as by wave breaking, it is accompanied by a shedding of momentum density to the underlying mean flow (Rapp & Melville, 1990). This connection between wave energy and mean momentum helps to illuminate the close two-way coupling between currents and waves, and as such will be of particular interest to the air-sea interaction community seeking to model transport and energy budgets between the atmosphere and ocean.
@@
In addition, we explored a connection between the mean potential energy, mean kinetic energy, and mean pressure, showing that at the surface, mean kinetic and mean potential energies are equal, which relates the surface drift directly to the mean water level. This is especially relevant to the observational community as direct measurements of the mean Lagrangian drift are particularly difficult (Kenyon, 1969), especially close to the surface (Lenain & Pizzo, 2020). On the other hand, measurements of geometric properties such the Lagrangian mean water level might offer an alternative way to estimate the mean Lagrangian drift, as in McAllister & van den Bremer (2019). One added caveat is that all of this analysis rests upon the assumption of two-dimensional flow, which is fundamentally different than fully three-dimensional flow, due to the fact that vorticity is no longer conserved on fluid particles because of vortex tilting/stretching. Examples of how to employ the Lagrangian frame for three-dimensional flow can be found in Yakubovich & Zenkovich (2001).
@@
It should be noted that the results presented here are similar to the “pseudomomentum rule” in Generalized Lagrangian Mean (GLM) theory, which states that $O(A^{2})$ mean forces can be calculated as if pseudomomentum were momentum, and the fluid medium were absent (McIntyre, 2019), where $A$ is the small wave amplitude. For surface waves, pseudomomentum per unit mass is defined as the wave energy over the phase speed $c$. We emphasize that it is twice the mean kinetic energy, which when divided by $c$, yields $U(\beta)$, valid to all orders of amplitude. To lowest order, the mean kinetic and potential energies are equal, which explains the $O(A^{2})$ result. The exact difference between mean kinetic and potential energy is given by the mean pressure (44), which has nonvanishing terms starting at $O(A^{4})$. The field of wave-mean interactions is vast (Bühler, 2014; Leibovich, 1983; Thomas, 2016), and while we do not investigate a general connection between momentum, energy, and vorticity for different types of waves, one may look towards Salmon (2016); Wagner & Young (2015); Thomas (2023) for other examples, a purely Lagrangian framework may prove insightful to such systems.
@@
Note that these results do not hold in the general case of rotational waves, such as Gerstner (1802) waves, which have no mean Lagrangian drift and therefore no net Lagrangian momentum density, yet still have a non-zero energy from their orbital motion. In this case, the connection between kinetic energy and drift fails due to the non-vanishing circulation. Here we focused solely on irrotational flow, though if desired any arbitrary vorticity could be prescribed to the system, generating a nonzero circulation which would balance the kinetic energy term in lieu of the drift. A comprehensive investigation of rotational flow is explored in the appendix.
@@
Acknowledgements. We thank the three anonymous referees whose constructive comments significantly improved the manuscript. We thank R. Salmon for helpful comments.
Funding. Blaser, Pizzo, and Lenain were partially supported by NSF OCE-2219752 and 2342714, NASA 80NSSC19K1037 (S-MODE) and 80NSSC23K0985 (OVWST) awards. Villas Bôas was supported by NASA 80NSSC23K0979 (OVWST) and NASA 80NSSC24K0411 (S-MODE) awards.
@@
Author Contributions.
Blaser
Formal analysis (lead); writing - original draft presentation (lead); writing - review and editing (lead); conceptualization (supporting). Benamran Formal analysis (supporting); writing - original draft presentation (supporting); conceptualization (supporting). Villas Bôas Conceptualization (supporting); writing - review and editing (supporting). Lenain Conceptualization (supporting); Funding Acquisition (lead for AB and RB); Supervision (supporting); Writing - review and editing (supporting); Pizzo Conceptualization (lead); Formal analysis (supporting); Writing - review and editing (supporting); Funding Acquisition (supporting); Supervision (lead).
Author ORCIDs.
[…]
$0$-0002-3104-4589 Aidan Blaser https://orcid.org/$0$-0002-3104-4589
[…]